7. Assume that a procedure yields a binomial distribution with a trial repeated n times. Use (l point) the binomial probability formula to find the probability ofx successes given the probability p of success on a single trial. Round to three decimal places. 0.015 0.012 0.004 0.023
Answer #1

Solution :

=> Answer :- Option a. 0.015

Given that n = 4 , x = 3 , p = 1/6 = 0.1667

=> q = 1 - p = 0.8333

=> P(x = 3) = 4C3*(0.1667^3)*(0.833^(4 - 3))

= (4!/3!*1!)*(0.1667^3)*(0.833^(4 - 3))

= 0.0154

= 0.015 (rounded)

Formula :

=> For binomial distribution , nCr*p^r*q^(n-r)

nCr = n!/(r!*(n-r)!)